OpenStudy (anonymous):

Can you simplify 1/18

OpenStudy (anonymous):

depends on what you mean by simplify it cannot be reduced

Parth (parthkohli):

Yes, \(\dfrac{0.5}{9}\), if that's what you call "simplify".

OpenStudy (anonymous):

one does not mix decimals and fractions

Parth (parthkohli):

That *is* why said, "if that's what you call 'simplify'" ;)

OpenStudy (dls):

no..but u can do the oppsite :) \[\LARGE \frac{(\sin^{2} \theta+\cos^{2} \theta)(1+w+w^{2})(\frac{7^{99}}{7^{98}}(\frac {7 \pi} {6})}{(\sqrt{100}+\sqrt{16}+2 ) \times 7(\frac{7 \pi}{6})} \]

OpenStudy (anonymous):

Is "no" a hard answer?

OpenStudy (anonymous):

If a fraction has a 1 in the numerator @shaneikque13 it is assumed it can no longer be simplified in basic math.